WellWheats, Inc. produces breakfast cereal and sells each box, or unit, for $7. The company has forecast production for the next three months as follows: July 5,000 units, August 6,000 units, September 3,500 units. Monthly manufacturing overhead is budgeted to be $20,000 plus $5 per unit produced. What is budgeted manufacturing overhead for July?

WellWheats, Inc. produces breakfast cereal and sells each box, or unit, for $7. The company is projecting sales of 1,000 units for the month of March. There are 30 units in the beginning inventory. Each unit requires 20 ounces of raw materials and 0.20 direct labor hours to make. The company's policy is to keep ending finished goods inventory of 10% of the current month's sales. Selling and administrative expenses for the month have been budgeted at $2,000.

Knowledge Check 02

Assume that the beginning raw materials inventory is 2,000 ounces. If the desired ending raw materials inventory is 20% of the current month's production needs, calculate the budgeted raw materials to be purchased.

Knowledge Check 03

If the direct labor cost per hour is $0.75, calculate the budgeted direct labor cost for the month of March.

WellWheats, Inc. produces breakfast cereal and sells each box, or unit, for $7. The company is projecting sales of 1,000 units for the month of March. There are 30 units in the beginning inventory. Each unit requires 20 ounces of raw materials and 0.20 direct labor hours to make. The company's policy is to keep ending finished goods inventory of 10% of the current month's sales. Selling and administrative expenses for the month have been budgeted at $2,000.

How many units should WellWheats produce in March?
Answer #1

Manufacturing overhead budget for July = Variable overhead + Fixed overhead

= 5,000*5 + 20,000

= $45,000

Estimated sales = 1,000 units

Add: Desired ending inventory = 100

Total required = 1,100 units

Less: beginning Inventory = 30

Production = 1,070 units

Raw material per unit = 20 Ounces

Total raw material for production = 21,400

Add: Desired ending inventory = 21,400*20% = 4,280

Less: beginning Inventory = 2,000

Budgeted Raw material to be purchased = 23,680 ounces

Production units = 1,070

Labor hours per unit = 0.20

Required labor hours = 214

Rate per hour = 0.75

Budgeted labor cost = 160.5

Estimated sales = 1,000 units

Add: Desired ending inventory = 100

Total required = 1,100 units

Less: beginning Inventory = 30

Production = 1,070 units